# LSEAttention is All You Need for Time Series Forecasting

### v5

@@ 2.2 Motivation
To address these challenges, We propose a module termed LSEAttention, which integrates the Log-Sum-Exp (LSE) defined as Eqn4 trick proposed by Blanchard et al. [10] along with the Gaussian Error linear Unit (GELU) activation function introduced by Hendrycks et al. [11]. The LSE trick alleviates entropy collapse issue arising from numerical instability through subtraction of max value. By starting from LSE where the softmax can be represented with LSE in this way:
$\text{softmax}(x_{i})=\frac{e^{x_{i}}}{\sum_{j=1}^{n}e^{x_{j}}}=e^{x_{i}-\text{LSE}(x)}.$ (5)
@@
$\frac{e^{x_{i}}}{e^{\log\left(\sum_{i=1}^{n}e^{x_{i}}\right)}}=\frac{e^{x_{i}}}{\sum_{j=1}^{n}e^{x_{j}}}=\text{softmax}(x_{i})$ (10)
where $g_{i}$ represents the $i$-th component of the stabilized softmax output.
@@ 3.1 Dataset
Sequence and Prediction Length: In accordance with prior works, We set the input sequence length at 336 time steps. The prediction length, which indicates the number of future time steps for which We aim to forecast data, varies among the set {96, 192, 336, 720}. This flexibility in prediction length allows for an evaluation of the model’s performance across different forecasting horizons.
@@ 3.2 Baseline
Hyperparameters: For all experiments, the hyperparameter settings for LATST are derived from those used in PatchTST, with a batch size fixed at 8 for simplicity.
Table 1: Hyperparameter settings used in experiments.
Hyperparameter Values
Batch size 8
Initial learning rate 0.5 to 0.00005
Number of heads 4 to 16
Head dropout rate 0, 0.2, 0.3
Input sequence length 336
Prediction length 96, 192, 336, 720
@@
Methods LATST SAMformer Transformer TSMixer iTrans Dlinear
ECL 96 0.131 0.155 0.182 0.173 0.148 0.140
192 0.154 0.168 0.202 0.204 0.162 0.153
336 0.169 0.183 0.212 0.217 0.178 0.169
720 0.197 0.219 0.238 0.242 0.225 0.203
Traffic 96 0.380 0.407 0.420 0.409 0.395 0.410
192 0.406 0.415 0.441 0.637 0.417 0.423
336 0.417 0.421 0.501 0.747 0.433 0.436
720 0.455 0.456 0.468 0.688 0.467 0.466
Weather 96 0.152 0.197 0.227 0.214 0.174 0.176
192 0.196 0.235 0.256 0.231 0.221 0.220
336 0.249 0.276 0.278 0.279 0.278 0.265
720 0.317 0.334 0.353 0.343 0.358 0.323
ETTh1 96 0.390 0.381 0.509 0.398 0.386 0.375
192 0.426 0.409 0.535 0.426 0.441 0.405
336 0.414 0.423 0.570 0.435 0.487 0.439
720 0.426 0.427 0.601 0.498 0.503 0.472
ETTh2 96 0.290 0.295 0.396 0.308 0.297 0.289
192 0.348 0.340 0.413 0.352 0.380 0.383
336 0.380 0.350 0.414 0.360 0.428 0.448
720 0.399 0.391 0.424 0.409 0.427 0.605
ETTm1 96 0.295 0.329 0.384 0.336 0.334 0.299
192 0.345 0.353 0.400 0.362 0.377 0.335
336 0.375 0.382 0.461 0.391 0.426 0.369
720 0.431 0.429 0.463 0.450 0.491 0.425
ETTm2 96 0.176 0.181 0.200 0.211 0.180 0.167
192 0.229 0.233 0.273 0.252 0.250 0.224
336 0.280 0.285 0.310 0.303 0.311 0.281
720 0.376 0.375 0.426 0.390 0.412 0.397
Exchange 96 0.098 0.161 0.292 0.343 0.086 0.081
192 0.218 0.246 0.372 0.342 0.177 0.157
336 0.408 0.368 0.494 0.484 0.331 0.305
720 0.865 1.003 1.323 1.204 0.847 0.603
Table 3: Parameters comparison of LATST and SAMformer models for various datasets and prediction lengths
Methods LATST SAMformer
Prediction Length 96 192 336 720 96 192 336 720
Electricity 251K 263K 282K 331K 82K 131K 205K 401K
Traffic 251K 263K 282K 331K 82K 131K 205K 401K
Weather 51K 54K 58K 72K 82K 131K 205K 401K
ETTH1 29K 29K 58K 331K 82K 131K 205K 401K
ETTH2 25K 27K 63K 331K 82K 131K 205K 401K
ETTM1 109K 115K 125K 331K 82K 131K 205K 401K
ETTM2 109K 115K 125K 150K 82K 131K 205K 401K
Exchange 51K 115K 199K 236K 82K 131K 205K 401K
@@ 3.3 Performance Analysis
The detailed performance metrics for each dataset and prediction length are summarized in Tab. 2. Notably, LATST consistently achieves the lowest MSE across various time-series datasets that exhibit substantial variability, which highlights its capacity to effectively exchange information between different variables. This phenomenon is closely related to the number of variables and parameter settings, where the model’s performance improves proportionally with the increasing complexity of interactions. A more in-depth analysis can be found in Tab. 3.

### v6

@@ 2.2 Motivation
To address these challenges, We propose a module termed LSEAttention, which integrates the Log-Sum-Exp (LSE) defined as Eq. 4 trick proposed by Blanchard et al. [10] along with the Gaussian Error linear Unit (GELU) activation function introduced by Hendrycks et al. [11]. The LSE trick alleviates entropy collapse issue arising from numerical instability through subtraction of max value. By starting from LSE where the softmax can be represented with LSE in this way:
$\text{softmax}(x_{i})=\frac{e^{x_{i}}}{\sum_{i=1}^{n}e^{x_{i}}}=e^{x_{i}-\text{LSE}(x)}.$ (5)
@@
$\frac{e^{x_{i}}}{e^{\log\left(\sum_{i=1}^{n}e^{x_{i}}\right)}}=\frac{e^{x_{i}}}{\sum_{i=1}^{n}e^{x_{i}}}=\text{softmax}(x_{i})$ (10)
where $g_{i}$ represents the $i$-th component of the stabilized softmax output. Instead of strictly converting back to original softmax function in Eq. 10. We instead of applying softmax to the $g_{i}$ to normalize the final attention weights into the bound between 0 and 1 in Eq. 11.
$\text{softmax}(g_{i})$ (11)
@@ 3.1 Dataset
Sequence and Prediction Length: In accordance with prior works, We set the input sequence length at 512 time steps for all models. The prediction length, which indicates the number of future time steps for which We aim to forecast data, varies among the set {96, 192, 336, 720}. This flexibility in prediction length allows for an evaluation of the model’s performance across different forecasting horizons.
@@ 3.2 Baseline
Hyperparameters: For all experiments, the hyperparameter settings for LATST are derived from those used in PatchTST, with a batch size fixed at 8 for simplicity. Rest of the hyparameters could be referred in Tab. 2 where it demonsteates clear pattern between the set of parameters and the categories.
@@
Methods LATST SAMformer PatchTST Transformer TSMixer iTrans Dlinear
ECL 96 0.130 0.155 - 0.182 0.173 - 0.140
192 0.151 0.168 - 0.202 0.204 - 0.153
336 0.169 0.183 - 0.212 0.217 - 0.169
720 0.197 0.219 - 0.238 0.242 - 0.203
Traffic 96 0.369 0.407 0.462 0.420 0.409 0.395 0.410
192 0.386 0.415 0.466 0.441 0.637 0.417 0.423
336 0.417 0.421 0.482 0.501 0.747 0.433 0.436
720 0.455 0.456 0.514 0.468 0.688 0.467 0.466
Weather 96 0.146 0.197 0.177 0.227 0.214 0.174 0.176
192 0.196 0.235 0.225 0.256 0.231 0.221 0.220
336 0.249 0.276 0.278 0.278 0.279 0.278 0.265
720 0.317 0.334 0.354 0.353 0.343 0.358 0.323
ETTh1 96 0.381 0.381 0.414 0.509 0.398 0.386 0.375
192 0.411 0.409 0.460 0.535 0.426 0.441 0.405
336 0.433 0.423 0.501 0.570 0.435 0.487 0.439
720 0.426 0.427 0.500 0.601 0.498 0.503 0.472
ETTh2 96 0.280 0.295 0.302 0.396 0.308 0.297 0.289
192 0.343 0.340 0.388 0.413 0.352 0.380 0.383
336 0.380 0.350 0.426 0.414 0.360 0.428 0.448
720 0.399 0.391 0.431 0.424 0.409 0.427 0.605
ETTm1 96 0.299 0.329 0.329 0.384 0.336 0.334 0.299
192 0.342 0.353 0.367 0.400 0.362 0.377 0.335
336 0.375 0.382 0.399 0.461 0.391 0.426 0.369
720 0.431 0.429 0.454 0.463 0.450 0.491 0.425
ETTm2 96 0.174 0.181 0.175 0.200 0.211 0.180 0.167
192 0.230 0.233 0.241 0.273 0.252 0.250 0.224
336 0.280 0.285 0.305 0.310 0.303 0.311 0.281
720 0.376 0.375 0.402 0.426 0.390 0.412 0.397
Exchange 96 0.102 0.161 0.088 0.292 0.343 0.086 0.081
192 0.211 0.246 0.176 0.372 0.342 0.177 0.157
336 0.408 0.368 0.301 0.494 0.484 0.331 0.305
720 0.865 1.003 0.901 1.323 1.204 0.847 0.603
Table 2: Classifications of Dataset where Low Variates: N = 4, D = 32, F = 64; Mid Variates: N = 8 D = 32, F = 128; High Variates: N = 16, D = 128, F = 256; Ultra-High Variates: N = 16 D = 256, F = 512. It demonstrates a clear pattern between number of variates and number of heads, as well as the dimension sizes.
Dataset Type
Low Variates ETTH1
ETTH2
ETTM1
ETTM2
Exchange
Mid Variates Weather
High Variates ECL
Ultra-High Variates Traffic
@@ 3.3 Performance Analysis
The detailed performance metrics for each dataset and prediction length are summarized in Tab. 1. Notably, LATST consistently achieves the lowest MSE across various time-series datasets that exhibit substantial variability, which highlights its capacity to effectively exchange information between different variables. This phenomenon is closely related to the number of variables and parameter settings, where the model’s performance improves proportionally with the increasing complexity of interactions. A more in-depth analysis can be found in LABEL:tab:parms.